# Multivariate Fields of Experts

### v1

@@ Abstract
We introduce the multivariate fields of experts, a new framework for the learning of image priors. Our model generalizes existing fields of experts methods by incorporating multivariate potential functions constructed via Moreau envelopes of the $\ell_{\infty}$-norm. We demonstrate the effectiveness of our proposal across a range of inverse problems that include image denoising, deblurring, compressed-sensing magnetic-resonance imaging, and computed tomography. The proposed approach outperforms comparable univariate models and achieves performance close to that of deep-learning-based regularizers while being significantly faster, requiring fewer parameters, and being trained on substantially fewer data. In addition, our model retains a relatively high level of interpretability due to its structured design.
Index Terms:
Learnable regularizer, variational methods, inverse problems, Moreau envelope.
@@ I Introduction
In many areas of science and engineering, it is common to recover an object of interest from indirect linear measurements. Let $\mathbf{H}\in\mathbb{R}^{m\times n}$ denote the measurement operator and let $\mathbf{y}\in\mathbb{R}^{m}$ be the observed data. The goal is to reconstruct the underlying signal $\mathbf{x}\in\mathbb{R}^{n}$ such that $\mathbf{H}\mathbf{x}\approx\mathbf{y}$. However, since $\mathbf{y}$ is usually noisy and $\mathbf{H}$ is often ill-conditioned or even rank-deficient, the direct inversion is typically unstable and results in poor reconstructions. To address this, a popular strategy is to use variational regularization to recover $\mathbf{x}$ as the minimizer of the energy
$\frac{1}{2}\|\mathbf{H}\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda R(\mathbf{x}),$ (1)
where the first term enforces data fidelity and where the regularizer $R\colon\mathbb{R}^{n}\to\mathbb{R}$ encodes prior knowledge about the underlying data.
Many popular regularizers [1, 2, 3, 4, 5, 6, 7] can be expressed as a sum of pointwise potential functions applied to the data convolved with a filter. In particular, the weakly convex ridge regularizer (WCRR) [5] represents a recent advance in this direction, where univariate nonconvex potentials are learned via linear splines.
In this work, we propose the multivariate fields of experts [^1] https://github.com/StanislasDucotterd/MFoE (MFoE), a novel class of learned regularizers that extends the classic fields of experts (FoE) framework by incorporating multivariate potential functions. Our contributions are as follows.
  • We introduce a class of parametric potentials that almost exactly reproduce the spline-based potentials learned in WCRR;
  • we extend WCRR to the multivariate setting by incorporating the $\ell_{\infty}$-norm into this parametric potential framework;
  • we demonstrate the effectiveness of our approach across multiple inverse problems: denoising, deblurring, compressed-sensing magnetic-resonance imaging (CS-MRI), and computed tomography (CT).
@@ II Related Work
A popular model for image reconstruction is the total variation (TV) regularizer [8], where the regularization term $R$ promotes piecewise-constant solutions by penalizing the horizontal and vertical finite differences of the image. Since $R$ is not differentiable, the objective (1) is typically minimized via the proximal operator
@@
$\operatorname{prox}_{\lambda R}(\mathbf{y})=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{n}}\left(\frac{1}{2}\|\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda R(\mathbf{x})\right),$ (2)
@@
which yields the solution of a denoising problem—the simplest instance of (1) with $\mathbf{H}=\mathbf{I}$. This proximal operator does not admit a closed-form solution and must be computed using iterative algorithms [9].
@@
To further improve the reconstruction performance of variational methods, many works have proposed data-driven approaches to construct $R$. The FoE framework [1] introduced a regularizer defined, for any image $\mathbf{x}\in\mathbb{R}^{n}$, as
$R(\mathbf{x})=\sum_{k=1}^{K}\langle\mathbf{1}_{n},\psi_{k}(\mathbf{W}_{k}\mathbf{x})\rangle,$ (3)
@@
where $\mathbf{W}_{k}\in\mathbb{R}^{n\times n}$ denotes a convolution matrix and $\psi_{k}\colon\mathbb{R}\to\mathbb{R}$ is a pointwise nonlinearity. In the original FoE model, the potentials $\psi_{k}$ correspond to the negative log-likelihood of Student-$t$ distributions and the filters $\{\mathbf{W}_{k}\}_{k=1}^{K}$ are learned by the minimization of $\sum_{m=1}^{M}R(\mathbf{x}_{m})$ over a dataset $\{\mathbf{x}_{m}\}_{m=1}^{M}$ of natural images .
@@
A more advanced strategy for the learning of the regularizer was proposed in [2], where bilevel optimization [10] is employed. In this setting, the filters are trained to minimize the denoising loss
$\sum_{m=1}^{M}\|\operatorname{prox}_{\lambda R}(\mathbf{x}_{m}+\mathbf{n}_{m})-\mathbf{x}_{m}\|_{2}^{2},$ (4)
@@
where $\mathbf{n}_{m}$ is additive white Gaussian noise. Although the potentials $\{\psi_{k}\}_{k=1}^{K}$ are generally nonconvex in FoE, which makes the proximal operator $\operatorname{prox}_{\lambda R}$ potentially set-valued, bilevel optimization remains applicable and effective in practice.
@@
Subsequent works [3, 4, 5, 6, 7] have further increased expressivity of the FoE model by jointly learning the filters and the nonlinearities $\{\psi_{k}\}_{k=1}^{K}$. While the WCRR learns the gradient of $R$ by parameterizing the derivatives of univariate potentials with linear splines, our model extends the approach to the multivariate setting using higher-dimensional potential functions.
@@
Another important framework for image reconstruction is plug-and-play (PnP) [11], where an off-the-shelf denoiser is inserted as a surrogate for the proximal operator in iterative optimization algorithms. While PnP methods are inspired by variational formulations, they generally do not correspond to the minimization of a well-defined energy functional and therefore lack convergence guarantees in general. Several strategies have been proposed to address this issue, such as Lipschitz constraints on the denoiser [12, 13], or continuation schemes [14] that gradually reduce the influence of the denoiser over the course of optimization.
@@
Finally, a recent line of work have bridged the gap between variational methods and the PnP framework by learning an explicit regularizer $R\colon\mathbb{R}^{n}\to\mathbb{R}$ using deep neural networks [15, 16, 17]. These methods supervise the model via its gradient $\nabla R$, which requires the computation of second-order derivatives during training. This yields highly expressive regularizers with an explicit objective and general convergence guarantees, but comes at the cost of significant computational overhead—both during training, due to the need to differentiate the objective twice, and at inference time, where the gradient of the neural network must be evaluated at each optimization step.
@@ III-A Notations and Definitions
The Moreau envelope [18] of a proper, lower semi-continuous, convex function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ with parameter $\mu>0$ is defined as
@@
The gradient of the Moreau envelope of $f$ can be related to the proximal operator of $f$ through the equality
@@
In this paper, we denote the Moreau envelope (with parameter $\mu$) of the $\ell_{\infty}$-norm in $\mathbb{R}^{d}$ as $\rho_{d}^{\mu}\colon\mathbb{R}^{d}\to\mathbb{R}_{+}$. Using the definition in (7) and the identity
$\operatorname{prox}_{\mu\|\cdot\|_{\infty}}(\mathbf{x})=\mathbf{x}-\mu\operatorname{Proj}_{\mathcal{B}_{\ell_{1}}}(\mathbf{x}/\mu)$ (9)
with $\mathbf{x}\in\mathbb{R}^{d}$, we can write that
@@
From (8), we then obtain the gradient
@@
which is computed via sorting using the Condat algorithm [19].
@@ III-B Generalization of the WCRR
In Figure 1, we illustrate that the potential learned via linear splines with more than 100 knots in the WCRR can be approximated almost to perfection by a simple parametric form with three degrees of freedom: Indeed, the function $t\mapsto\lambda(\rho_{1}^{\mu}(t)-\rho_{1}^{\nu}(t))$ with $\mu=0.002$ and $\lambda=\nu=0.098$ is visually indistinguishable from the WCRR of Figure 1. This type of potential is close to the minimax-concave penalty introduced in [20]. This observation motivates the use of $\rho_{d}^{\mu}$ as a foundation for a multivariate extension, with the property that the model can reduce to the WCRR when $d=1$.
Note that the absolute value is the only one-dimensional function (up to scaling) that is a norm. This suggests to extend it by the consideration of the Moreau envelopes of general norms. We focus on norms for which both the Moreau envelope and its gradient can be computed efficiently. Among the classic choices, the $\ell_{2}$ and $\ell_{\infty}$ norms are particularly tractable. We opt for the $\ell_{\infty}$ norm for two main reasons: first, it empirically leads to better reconstruction results; second, as established in [21], any norm can be approximated arbitrarily well by the composition of the $\ell_{\infty}$ norm with a linear transformation. This makes it a particularly expressive and computationally attractive choice for multivariate potentials.
The WCRR constrains its potentials to be 1-weakly convex to ensure overall convexity during denoising, given the 1-strong convexity of the data term. While this has minimal impact in the univariate case, it limits performance in the multivariate setting. We therefore relax this constraint in our model and allow potentials with larger weak-convexity constants. To isolate the effect of our multivariate formulation from that of the relaxation, we also compare against a relaxed version of WCRR.
Figure 1: Top: Potential learned in WCRR, vertical lines mark the spline control points. Bottom: Difference between two Moreau envelopes of the absolute value with different parameters.
@@ III-C General Setting
We now move beyond the one-dimensional setting and extend the FoE framework to the multivariate case, where the regularizer is defined, for any image $\mathbf{x}\in\mathbb{R}^{n}$, as
@@
where $\mathbf{W}_{k}^{d}\in\mathbb{R}^{(dn)\times n}$ denotes a convolution matrix that maps a single-channel image to a $d$-channel one, and $\psi_{k}^{d}\colon\mathbb{R}^{d}\to\mathbb{R}$ is a nonlinearity given by
$\psi_{k}(\mathbf{x})=\mu_{k}\rho_{\mu_{k}}^{d}(\mathbf{x})-\mu_{k}\rho_{\tau_{k}\mu_{k}}^{d}(\mathbf{Q}_{k}\mathbf{x}),$ (13)
@@
If $\mathbf{Q}_{k}$ is such that $\|\mathbf{Q}_{k}\|_{\infty}\leq 1$ and $\tau_{k}>\|\mathbf{Q}_{k}\|_{2}^{2}$, then the nonlinearity defined in (13) is nonnegative with a unique global minimum at the origin and has nonexpansive gradients.
Theorem 1 guarantees that the regularizer is bounded from below and preserves a similar interpretation to the classic FoE, where weak filter responses are favored while stronger responses are increasingly penalized. The proof can be found in Appendix -A.
@@
where $\mathbf{W}\in\mathbb{R}^{(Kdn)\times n}$ refers to the concatenation of all the $d$-channel convolutions $\mathbf{W}_{k}^{d}$ and $\varphi$ refers to the concatenation of all the gradients $\mathbf{\nabla}\psi_{k}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$. The spectral norm of the filters $\|\mathbf{W}\|_{2}$ is enforced to be one via the same methods as in [5], which ensures that $\mathbf{\nabla}R$ is nonexpansive.
We train the filters $\mathbf{W}$, the $K$ matrices $\mathbf{Q}_{k}$, scalars $\mu_{k},\tau_{k}$, and the regularizing strength $\lambda$ to minimize the denoising loss (4).
To efficiently train filters with a large receptive field, we decompose $\mathbf{W}$ into a composition of three zero-padded convolutions with small kernels and an increasing number of output channels as in [4]. Similarly to [22], the filters are constrained to have zero mean.
@@ III-D Optimization
We minimize the objective (1) using accelerated gradient descent with restart based on the value of the objective as described in [23]. We give the pseudocode for our implementation in Algorithm 1 where $f$ corresponds to (1). Throughout the paper, we set
$\gamma=\frac{1}{\|\mathbf{H}\|_{2}^{2}+\lambda}.$ (15)
Since the regularizer has a gradient $\mathbf{\nabla}R$ that is nonexpansive, the Lipschitz constant of the gradient of (1) is upper-bounded by the Lipschitz constant of the data-fitting term plus the regularization parameter $\lambda$.
The computational cost of the fourth line of Algorithm 1 turns out to be negligible, due to the nature of our potentials. Indeed, from (10) and (11), we have that
$\rho_{d}^{\mu}(\mathbf{x})=\|\mathbf{x}-\mu\nabla\rho_{d}^{\mu}(\mathbf{x})\|_{\infty}+\frac{\mu}{2}\|\nabla\rho_{d}^{\mu}(\mathbf{x})\|_{2}^{2},$ (16)
which makes it very easy to compute the objective on top of the gradient. Therefore, the restart strategy allows us to decrease the amount of iterations without increasing the time required to do one.
Algorithm 1 Accelerated gradient descent with restart based on the objective
@@
1: Input: Function $f$, stepsize $\gamma$, initialization $\mathbf{x}_{0}\in\mathbb{R}^{d}$, tolerance $\varepsilon>0$, number of iterations $n_{\text{max}}$
2:Set $t_{0}=1$, $k=0$, $r_{0}=\infty$, $f_{0}=\infty$, $\mathbf{z}_{0}=\mathbf{x}_{0}$
3:while $r_{k}>\varepsilon$ and $k\leq n_{\text{max}}$ do
4:  $\mathbf{x}_{k+1}=\mathbf{z}_{k}-\gamma\nabla f(\mathbf{z}_{k})$
5:  $f_{k+1}=f(\mathbf{z}_{k})$
6:  $t_{k+1}=\frac{1+\sqrt{1+4t_{k}^{2}}}{2}$
7:  $\mathbf{z}_{k+1}=\mathbf{x}_{k+1}+\frac{t_{k}-1}{t_{k+1}}(\mathbf{x}_{k+1}-\mathbf{x}_{k})$
8:  if $f_{k+1}>f_{k}$ then
9:   $\mathbf{z}_{k+1}=\mathbf{x}_{k+1}$
10:   $t_{k+1}=1$
11:  end if
12:  $r_{k+1}=\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\|/\|\mathbf{x}_{k+1}\|$
13:  $k\leftarrow k+1$
14:end while
15: Output: Approximate solution $\mathbf{x}_{k+1}$
III-E Dependence on the Noise Level
To make our regularizer more universal, we train it on a range of Gaussian-noise levels and denote by $R_{\sigma}$ the regularizer associated with the noise level $\sigma$. To keep the model as simple as possible, we share the filters and most other parameters across noise levels, allowing only $\{\mu_{k}\}_{k=1}^{K}$ to depend on $\sigma$. We define a function $\boldsymbol{\mu}\colon\mathbb{R}\to\mathbb{R}^{K}$ that concatenates the $\mu_{k}$, such that $\mu_{k}(\sigma)=\mu(\sigma)_{k}$. This function is parameterized as
$\boldsymbol{\mu}(\sigma)=\frac{1}{100}\left(\sigma\cdot\mathbf{1}_{K}+\frac{1}{20}\mathbf{f}(\bar{\sigma})\right)_{+}+10^{-9}\cdot\mathbf{1}_{K},$ (17)
where $\mathbf{f}\colon\mathbb{R}\to\mathbb{R}^{K}$ is a three-layer neural network. The function $(\cdot)_{+}\coloneqq\max(0,\cdot)$ and the addition of a small term are here to ensure that components of $\boldsymbol{\mu}$ stay positive. The term $\bar{\sigma}$ refers to a min-max normalization of $\sigma$ such that it is in the interval [2,2][-2,2] in order to fully benefit from the neural network $\mathbf{f}$ whose knots at initialization are mostly located in that range.
IV Experiments
IV-A Training on Denoising
We denote by $\boldsymbol{\theta}$ the vector of all parameters in our model. The paired data used for training are $(\mathbf{x}_{m},\mathbf{y}_{m})_{m=1}^{M}$, where each noisy observation is generated as $\mathbf{y}_{m}=\mathbf{x}_{m}+\sigma_{m}\mathbf{n}_{m}$ with Gaussian noise $\mathbf{n}_{m}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and noise levels $\sigma_{m}\sim\mathcal{U}(0,0.2)$. The model is trained to minimize the reconstruction loss

### v2

@@ Abstract
We introduce the multivariate fields of experts, a new framework for the learning of image priors. Our model generalizes existing fields of experts methods by incorporating multivariate potential functions constructed via Moreau envelopes of the $\ell_{\infty}$-norm. We demonstrate the effectiveness of our proposal across a range of inverse problems that include image denoising, deblurring, compressed-sensing magnetic-resonance imaging, and computed tomography. The proposed approach outperforms comparable univariate models and achieves performance close to that of deep-learning-based regularizers while being significantly faster, requiring fewer parameters, and being trained on substantially fewer data. In addition, our model retains a high level of interpretability due to its structured design. It is supported by theoretical convergence guarantees which ensure reliability in sensitive reconstruction tasks.
@@ I Introduction
A common problem in science and engineering is to recover an object of interest from indirect linear measurements. Let $\mathbf{H}\in\mathbb{R}^{m\times n}$ denote the measurement operator and let $\mathbf{y}\in\mathbb{R}^{m}$ be the observed data. The goal is to reconstruct the underlying signal $\mathbf{x}\in\mathbb{R}^{n}$ such that $\mathbf{H}\mathbf{x}\approx\mathbf{y}$. However, since $\mathbf{y}$ is usually noisy and $\mathbf{H}$ is often ill-conditioned or even rank-deficient, the direct inversion is typically unstable and results in poor reconstructions. To address this, a popular strategy is to use variational regularization to recover $\mathbf{x}$ as the minimizer of the energy
$f(\mathbf{x})=\tfrac{1}{2}\|\mathbf{H}\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda R(\mathbf{x}),$ (1)
where the first term enforces data fidelity and the regularizer $R\colon\mathbb{R}^{n}\to\mathbb{R}$ encodes prior knowledge about the underlying signal.
Many popular regularizers [34, 10, 27, 18, 19, 45, 32] can be expressed as a sum of univariate potential functions applied to filter responses, a structure frequently referred to as the fields of experts (FoE) model. In particular, the weakly convex ridge regularizer (WCRR) [19] represents a recent advance in this direction, where nonconvex potentials are learned via linear splines. While highly effective, such approaches implicitly assume independence between the filter responses (channels), ignoring potentially valuable interactions between them.
In this work, we propose the multivariate fields of experts (MFoE), a novel class of learned regularizers that generalizes the classic FoE framework by incorporating multivariate potential functions that have the capability to capture interactions between channels. Our contributions are as follows.
  • Multivariate Generalization: We extend the WCRR framework to the multivariate setting by introducing a novel class of parametric potentials based on Moreau envelopes.
  • Tailored Optimization: We design a specialized optimization algorithm for the resulting objective and prove its convergence to a stationary point.
  • Comprehensive Validation: We demonstrate the effectiveness of our approach across diverse inverse problems that include denoising, deblurring, compressed-sensing magnetic-resonance imaging (CS-MRI), and computed tomography (CT). We demonstrate superior performance to univariate FoE methods. Our code is accessible on Github  [^1] https://github.com/StanislasDucotterd/MFoE.
@@ II Related Work
A popular model for image reconstruction is the total variation (TV) regularizer [35], where the regularization term $R_{\text{TV}}$ promotes piecewise-constant solutions by penalizing the horizontal and vertical finite differences of the image. Since $R_{\text{TV}}$ is not differentiable, the objective (1) is typically minimized via the proximal operator
@@
$\operatorname{prox}_{\lambda R}(\mathbf{y})=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{n}}\left(\tfrac{1}{2}\|\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda R_{\text{TV}}(\mathbf{x})\right),$ (2)
@@
which yields the solution of a denoising problem—the simplest instance of (1) with $\mathbf{H}=\mathbf{I}$. This proximal operator does not admit a closed-form solution and must be computed through iterative algorithms [8].
@@
To further improve the reconstruction performance of variational methods, many works propose data-driven approaches to construct $R$. The FoE framework [34] relies on a regularizer defined for any image $\mathbf{x}\in\mathbb{R}^{n}$ as
$R_{\text{FoE}}(\mathbf{x})=\sum_{k=1}^{K}\langle\mathbf{1}_{n},\psi_{k}(\mathbf{W}_{k}\mathbf{x})\rangle,$ (3)
@@
where $\mathbf{W}_{k}\in\mathbb{R}^{n\times n}$ denotes a convolution matrix and $\psi_{k}\colon\mathbb{R}\to\mathbb{R}$ is a univariate nonlinearity. In the original FoE model, the potentials $\psi_{k}$ correspond to the negative log-likelihood of Student-$t$ distributions and the filters $\{\mathbf{W}_{k}\}_{k=1}^{K}$ are learned by the minimization of $\sum_{m=1}^{M}R_{\text{FoE}}(\mathbf{x}_{m})$ over a dataset $\{\mathbf{x}_{m}\}_{m=1}^{M}$ of natural images.
@@
A more advanced strategy for the learning of the regularizer is proposed in [10], where bilevel optimization [39] is employed. In this setting, the filters are trained explicitly to minimize the denoising loss
$\sum_{m=1}^{M}\|\operatorname{prox}_{\lambda R_{\text{FoE}}}(\mathbf{x}_{m}+\mathbf{n}_{m})-\mathbf{x}_{m}\|_{2}^{2},$ (4)
where $\mathbf{n}_{m}$ is an additive white Gaussian noise. This method aligns the training objective with the evaluation metric (PSNR), which yields better results than the likelihood-based learning of the original FoE.
@@
Although the potentials $\{\psi_{k}\}_{k=1}^{K}$ are generally nonconvex in FoE, which makes the proximal operator $\operatorname{prox}_{\lambda R}$ potentially set-valued, bilevel optimization remains applicable and effective in practice.
@@
Subsequent works [27, 18, 19, 45, 32] increase the expressivity of the FoE model further by jointly learning the filters and the nonlinearities $\{\psi_{k}\}_{k=1}^{K}$. The WCRR learns the gradient of the regularizer by parameterizing the derivatives of univariate potentials with linear splines. Our model relies on higher-dimensional potential functions to extend this approach to the multivariate setting.
@@
While the approaches discussed above minimize an explicit energy functional, a parallel line of research known as Plug-and-Play (PnP) [43] bypasses the use of an explicit regularizer. Instead, PnP methods insert an off-the-shelf denoiser as a surrogate for the proximal operator in iterative optimization algorithms. Although PnP yields good performance, it generally does not correspond to the minimization of an explicit objective and often lacks convergence guarantees. Several strategies have been proposed to address this issue, such as the enforcement of Lipschitz constraints on the denoiser [36, 13] or the use of continuation schemes [9].
@@
Finally, a recent line of work has bridged the gap between variational methods and the PnP framework by parameterizing an explicit regularizer with deep neural networks [20, 21, 15]. These methods learn the regularizer via its gradient, which requires the computation of second-order derivatives during training. This yields highly expressive regularizers with an explicit objective and general convergence guarantees, but comes at the cost of a significant computational overhead—both during training, due to the need to differentiate the objective twice, and at inference time, since the gradient of the neural network must be evaluated at each optimization step.
@@ III-A Notations and Definitions
The Moreau envelope [26] of a proper, lower semicontinuous, convex function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ with parameter $\mu>0$ is defined as
@@
The gradient of the Moreau envelope of $f$ is related to the proximal operator of $f$ through the equality
@@
In this paper, we denote the Moreau envelope (with parameter $\mu$) of the $\ell_{\infty}$-norm in $\mathbb{R}^{d}$ as $\rho^{d}_{\mu}\colon\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$. We rely on the standard identity [30, Equation 6.8]
$\operatorname{prox}_{\mu\|\cdot\|_{\infty}}(\mathbf{x})=\mathbf{x}-\mu\operatorname{Proj}_{\mathcal{B}_{\ell_{1}}}(\mathbf{x}/\mu).$ (9)
The proximal operator explicitly realizes the infimum in the definition of the Moreau envelope. Therefore, the evaluation of (7) at $\mathbf{z}=\operatorname{prox}_{\mu\|\cdot\|_{\infty}}(\mathbf{x})$ yields
@@
From (8), we then obtain the convenient formula for the gradient
@@
which can be computed efficiently via sorting using the Condat algorithm [11].
@@ III-B Generalization of the WCRR
The WCRR [19] constructs a regularization functional by applying 1-weakly convex, univariate potential functions parameterized as linear splines to learned filter responses. Our proposed model generalizes this framework by replacing these scalar functions with higher-dimensional analogs, a generalization directly motivated by the geometric properties of the WCRR potentials. Specifically, these potentials take the form $\psi_{k}(x)=\alpha_{k}^{-2}\psi(\alpha_{k}x)$ where a single base function $\psi$ is shared across all channels. In Figure 1, we illustrate that $\psi$ (which is learned via a linear spline with more than 100 knots) can be approximated almost to perfection by a simple three-parameter function of the form $t\mapsto\lambda(\rho_{\mu}^{1}(t)-\rho_{\nu}^{1}(t))$ with $\mu=0.002$ and $\lambda=\nu=0.098$. This observation motivates the use of $\rho^{d}_{\mu}$ as a foundation for a multivariate extension, with the property that the model can reduce to the WCRR when $d=1$. This three-parameter function is closely related to the minimax-concave penalty introduced in [37].
Note that the absolute value is the only one-dimensional function (up to scaling) that is a norm. This suggests to extend it by the consideration of the Moreau envelopes of general norms. We focus on norms for which both the Moreau envelope and its gradient can be computed efficiently. The $\ell_{1}$-norm decomposes into a sum of absolute values, which implies that the difference of two Moreau envelopes recovers the univariate formulation of the WCRR.
The $\ell_{2}$ and $\ell_{\infty}$ norms are particularly tractable. We select the $\ell_{\infty}$-norm for two reasons: first, it yields better empirical results; second, the authors of [40, Theorem 4] establish that any norm on $\mathbb{R}^{d}$ can be approximated to arbitrary precision by the function $\mathbf{x}\mapsto\|\mathbf{L}\mathbf{x}\|_{\infty}$, where $\mathbf{L}$ maps into a potentially higher-dimensional space. This universality property ensures that our model can represent a broad class of convex potentials simply by learning the appropriate linear filters.
The WCRR constrains its potentials to be 1-weakly convex to ensure overall convexity during denoising, given the 1-strong convexity of the data term. Our optimization algorithm allows us to relax this constraint in our model without losing any convergence guarantees.
Figure 1: The WCRR potential (top) and its gradient (bottom) match the difference of two Moreau envelopes of the absolute value. Vertical ticks indicate the spline knots.
@@ III-C General Setting
We now move beyond the univariate setting and extend the FoE framework to the multivariate case, where the regularizer is defined for any image $\mathbf{x}\in\mathbb{R}^{n}$ as
@@
where $\mathbf{W}_{k}^{d}\in\mathbb{R}^{(dn)\times n}$ denotes a multi-convolution matrix that maps a single-channel image to a $d$-channel one, and $\psi_{k}^{d}\colon\mathbb{R}^{d}\to\mathbb{R}$ is a nonlinearity given by
$\psi_{k}^{d}(\mathbf{x})=\mu_{k}\rho_{\mu_{k}}^{d}(\mathbf{x})-\mu_{k}\rho_{\tau_{k}\mu_{k}}^{d}(\mathbf{Q}_{k}\mathbf{x}),$ (13)
@@
If $\mathbf{Q}_{k}$ is such that $\|\mathbf{Q}_{k}\|_{\infty}\leq 1$ and $\tau_{k}>\|\mathbf{Q}_{k}\|_{2}^{2}$, then the nonlinearity (13) is nonnegative with a unique global minimum at the origin, and a gradient that is nonexpansive.
Theorem 1 guarantees that the regularizer is bounded from below and preserves a interpretation similar to the classic FoE, where weak filter responses are favored while stronger responses are penalized. The proof can be found in Appendix -A.
@@
where $\mathbf{W}\in\mathbb{R}^{(Kdn)\times n}$ refers to the concatenation of all the $d$-channel convolutions and $\varphi$ refers to the concatenation of all the gradients $\mathbf{\nabla}\psi_{k}^{d}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$. The spectral norm $\|\mathbf{W}\|_{2}$ of the filters is enforced to be one which ensures that $\mathbf{\nabla}R$ is nonexpansive.
To efficiently train filters with a large receptive field, we decompose $\mathbf{W}$ into a three zero-padded convolutions with small kernels and an increasing number of output channels, as in [18]. Similarly to [22], the filters are constrained to have zero mean.
@@ III-D Optimization
We design an algorithm that minimizes the objective function defined in (1) and guarantees convergence to a stationary point. We build upon the work of [28] which uses the heavy-ball method [31]
$\mathbf{x}_{k+1}=\mathbf{x}_{k}-\alpha_{k}\nabla f(\mathbf{x}_{k})+\beta_{k}(\mathbf{x}_{k}-\mathbf{x}_{k-1}).$ (15)
The method of [28] ensures convergence by imposing strict constraints on the stepsize $\alpha_{k}$ and momentum $\beta_{k}$. Here, we relax these constraints to allow for more aggressive acceleration while still maintaining convergence guarantees, with the backing of Theorem 2 below. As described in Algorithm 1 (Lines 5–6), our refinement is a backtracking mechanism that discards the inertial steps and falls back to a standard gradient descent update if the candidate update $\mathbf{x}_{k+1}$ fails to satisfy a sufficient descent condition.
The backtracking step introduces negligible overhead because the cost of the check of the descent condition is minimal. This is due to the specific structure of our potentials
$\rho^{d}_{\mu}(\mathbf{x})=\|\mathbf{x}-\mu\nabla\rho^{d}_{\mu}(\mathbf{x})\|_{\infty}+\frac{\mu}{2}\|\nabla\rho^{d}_{\mu}(\mathbf{x})\|_{2}^{2}.$ (16)
Since $R$ is composed of these potentials, the function value $R(\mathbf{x})$ can be computed cheaply alongside the gradient $\nabla R(\mathbf{x})$. Furthermore, the values computed during a rejected candidate step are reused for the fallback step, which ensures that the objective and gradient are evaluated twice only when a restart occurs.
The nonexpansiveness of $\mathbf{\nabla}R$ implies that $\mathbf{\nabla}f$ is $L$-Lipschitz continuous with constant
$L=\|\mathbf{H}\|_{2}^{2}+\lambda.$ (17)
Theorem 2
Let $\alpha\in(0,2/L)$. If the sequence $\{\mathbf{x}_{k}\}_{k\geq 0}$ generated by Algorithm 1 is bounded, then it converges to a stationary point of $f$ and has a finite length, with
$\sum_{k\geq 0}\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\|_{2}<\infty.$ (18)
Theorem 2 is essential to the reliability of image reconstruction, as the finite-length property guarantees that the iterates settle to a specific solution rather than being indefinitely oscillating or diverging. The proof can be found in Appendix -B. The theorem assumes bounded iterates, a condition that can also be satisfied formally by augmenting the objective with a coercive barrier term such as $B(\mathbf{x})=\log\left(1+\exp\left(\|\mathbf{x}\|_{2}^{2}-C\right)\right)$. There, the constant term $C$ can be chosen arbitrarily large. The coercive term ensures bounded level sets while it remains negligible within the valid range of image intensities. Similar ideas have been proposed in [19] and [21].
Throughout this work, we use the fixed stepsize $\alpha=1.99/L$. We increase the momentum parameter $\beta_{k}$ after a successful inertial step and reset it to a base value otherwise, a common heuristic in accelerated optimization.
Algorithm 1 Heavy-Ball with Restart
@@
1:Input: Function $f$ with $L$-Lipschitz gradients, stepsize $\alpha$, initialization $\mathbf{x}_{0}\in\mathbb{R}^{n}$, tolerance $\varepsilon>0$
2:Set $k=0$, $\beta_{0}=0.5$, $r=\infty$
3:while $r>\varepsilon$ do
4:  $\mathbf{x}_{k+1}=\mathbf{x}_{k}-\alpha\mathbf{\nabla}f(\mathbf{x}_{k})+\beta_{k}\left(\mathbf{x}_{k}-\mathbf{x}_{k-1}\right)$
5:  $\beta_{k+1}=\tfrac{1}{2}(\beta_{k}+1)$
6:  if $f(\mathbf{x}_{k+1})>f(\mathbf{x}_{k})-\alpha\left(1-\frac{L\alpha}{2}\right)\|\nabla f(\mathbf{x}_{k})\|_{2}^{2}$ then
7:   $\mathbf{x}_{k+1}=\mathbf{x}_{k}-\alpha\mathbf{\nabla}f(\mathbf{x}_{k})$
8:   $\beta_{k+1}=0.5$
9:  end if
10:  $r=\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\|/\|\mathbf{x}_{k+1}\|$
11:  $k\leftarrow k+1$
12:end while
13: Output: Approximate solution $\mathbf{x}_{k+1}$
IV Training on Denoising
We adopt a bilevel optimization strategy to learn the parameters
$\boldsymbol{\theta}=\{\mathbf{W},\{\mathbf{Q}_{k},\tau_{k}\}_{k=1}^{K},\boldsymbol{\theta}_{\mathbf{f}},\lambda\},$ (19)
where $\boldsymbol{\theta}_{\mathbf{f}}$ parametrizes the function $\mathbf{f}$ from Section -C. To enforce the constraints $\|\mathbf{W}\|_{2}=1$, $\|\mathbf{Q}_{k}\|_{\infty}\leq 1$, and $\tau_{k}>\|\mathbf{Q}_{k}\|_{2}^{2}$, we optimize over unconstrained auxiliary variables mapped to the feasible set via
$\mathbf{W}$ $=\frac{\tilde{\mathbf{W}}}{\|\tilde{\mathbf{W}}\|_{2}},\quad\mathbf{Q}_{k}=\frac{\tilde{\mathbf{Q}}_{k}}{\max(1,\|\tilde{\mathbf{Q}}_{k}\|_{\infty})},$ (20)
[…]
where the spectral norm $\|\tilde{\mathbf{W}}\|_{2}$ is computed via the Fourier transform [19].
IV-A Bilevel Formulation
To make our regularizer more universal, we train it on a range of Gaussian-noise levels and denote by $R_{\sigma}$ the regularizer associated with the noise level $\sigma$. We use a dataset of paired samples $(\mathbf{x}_{m},\mathbf{y}_{m})_{m=1}^{M}$, where $\mathbf{x}_{m}$ is the reference image and $\mathbf{y}_{m}=\mathbf{x}_{m}+\sigma_{m}\mathbf{n}_{m}$ with $\mathbf{n}_{m}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\sigma_{m}\sim\mathcal{U}(0,0.2)$. This yields the bilevel optimization problem